Listed below are the amounts of net worth​ (in millions of​ dollars) of the ten wealthiest celebrities in a country. Construct a

99​%

confidence interval. What does the result tell us about the population of all​ celebrities? Do the data appear to be from a normally distributed population as​ required?

266

192

190

167

159

154

154

154

152

149

What is the confidence interval estimate of the population mean

muμ​?

​$nothing

millionless than<muμless than<​$nothing

million

​(Round to one decimal place as​ needed.)
Answer #1

Solution:

Confidence interval = Xbar ± t*S/sqrt(n)

From given data, we have

Xbar = 173.7

S = 35.94146476

n = 10

df = n – 1 = 10 – 1 = 9

confidence level = 99%

Critical t value = 3.2498

(by using t-table)

Confidence interval = 173.7 ± 3.2498*35.94146476/sqrt(10)

Confidence interval = 173.7 ± 3.2498* 11.36568911

Confidence interval = 173.7 ± 36.9366

Lower limit = 173.7 - 36.9366 = 136.76

Upper limit = 173.7 + 36.9366 = 210.64

$136.8 million < µ < $210.6 million